Maximize Profit

A TV retailer supposes that in order to sell n nimber of TVs, the price per unit must follow the model p=600 - 0.3n. The retailer also supposes that the total cost of keeping n number of TVs in inventory is given by the model C(n) = 0.3n^2 + 5000. How many TVs must the retailer keep in inventory and sell in order to maximize his profit?
Answer #1

Profit = (600-0.3n) *n - (0.3n^2+5000)

= 600n-0.3n^2-0.3n^2-5000

=-0.6n^2+600n-5000

d(profit)/dn = -1.2n + 600 = 0

               n = 500


answered by: mathboy